### 01 Apr 2026 v3

# EagleNet: Energy-Aware Fine-Grained Relationship Learning Network for Text-Video Retrieval

Abstract
Text-video retrieval tasks have seen significant improvements due to the recent development of large-scale vision-language pre-trained models. Traditional methods primarily focus on video representations or cross-modal alignment, while recent works shift toward enriching text expressiveness to better match the rich semantics in videos. However, these methods use only interactions between text and frames/video, and ignore rich interactions among the internal frames within a video, so the final expanded text cannot capture frame contextual information, leading to disparities between text and video. In response, we introduce Energy-Aware Fine-Grained Relationship Learning Network (EagleNet) to generate accurate and context-aware enriched text embeddings. Specifically, the proposed Fine-Grained Relationship Learning mechanism (FRL) first constructs a text-frame graph by the generated text candidates and frames, then learns relationships among texts and frames, which are finally used to aggregate text candidates into an enriched text embedding that incorporates frame contextual information. To further improve fine-grained relationship learning in FRL, we design Energy-Aware Matching (EAM) to model the energy of text-frame interactions and thus accurately capture the distribution of real text-video pairs. Moreover, for more effective cross-modal alignment and stable training, we replace the conventional softmax-based contrastive loss with the sigmoid loss. Extensive experiments have demonstrated the superiority of EagleNet across MSRVTT, DiDeMo, MSVD, and VATEX. Codes are available at https://github.com/draym28/EagleNet.
1 Introduction
Figure 1: The performance-efficiency comparison among SOTA models (ViT-B/16) on MSRVTT datasets.
As the volume of video data being generated and stored continues to grow, the need for effective content retrieval models has become more urgent than ever. Text-Video Retrieval (TVR) has been extensively studied and shows great progress in recent years. In Retrieval-Augmented Generation (RAG) [15], TVR improves the accuracy of generative models by integrating video content into knowledge-based systems, enhancing user interactions in applications like dialogue systems.
Early works [55, 9, 49, 10, 58, 16, 14, 29, 26, 46, 5, 31, 13] mainly rely on pre-extracted features, which are learned by independent text/image encoders. Additionally, due to the huge development in Graph Neural Network (GNN), many researchers utilize graph learning techniques in text-video retrieval tasks. As the groundbreaker, HGR [5] constructs a semantic graph for each sentence, and utilizes GNN to learn text embeddings. Inspired by HGR, many follow-up works [44, 23, 19, 13] employ GNN for text-video structural relationship learning. For example, ACP [13] introduces knowledge graph learning model to help complete semantic information for retrieval. However, all of these works are based on data pre-processing approaches with heavy external knowledge or outdated encoders, limiting the flexibility and increasing the complexity of the model.
Instead, recent studies incorporate the emergent large-scale image-text pre-trained models into text-video retrieval tasks [33, 17, 34, 53, 48, 27, 43, 12, 32, 7, 38, 51, 54, 57, 3, 41]. They utilize CLIP image and text encoder [37] as multi-modal encoders to achieve strong performance on many text-video retrieval benchmarks. CLIP4Clip [33], the first work that applies CLIP in this task, proposes using a temporal aggregator to fuse frames and then align text and video features. Later, many works [17, 12, 32, 7, 57, 3, 41] focus on obtaining high-quality video features to enhance cross-modal matching. Besides, some other works [59, 27, 48, 54, 57, 38] propose to improve the text-video alignment, such as conducting matching at different granularities. However, most models overlook the insufficient expressiveness of text embeddings. The brevity of video descriptions often prevents text embeddings from fully reflecting the underlying video semantics. To fill this gap, TMASS [43] stochastically samples enriched text within a flexible range, the radius of which is determined by the similarity of text and frame embeddings. Similarly, TV-ProxyNet [51] refines the text query into a specific text proxy for each candidate video proxy through video-aware directors and text-video similarities. Nevertheless, these methods focus solely on interactions between text and frames/video when expanding text semantics, and ignore relationships among video frames, so the final enriched text cannot capture frame contextual information, leading to disparities between text and video.
We argue that the internal frame-to-frame relations within the video should also be incorporated when expanding text semantics. This enables the model to capture both global and temporal semantics of videos, allowing the generated text embeddings to reflect not only the semantics of individual frames but also their contextual dependencies. Moreover, by establishing relationships among video semantic units (frames), the model can effectively identify and suppress redundant information and noise, leading to more accurate, contextually enriched text embeddings.
To this end, we introduce the Energy-Aware Fine-Grained Relationship Learning Network (EagleNet), a framework designed to generate accurate and context-aware enriched text embeddings by exploring fine-grained relationships among text and video frames (see Fig. 2). The model consists of two key components: 1) Fine-Grained Relationship Learning (FRL) and 2) Energy-Aware Matching (EAM). Specifically, FRL first generates multiple text candidates using a stochastic text modeling strategy, and constructs a text-frame relational graph. We then design an effective relational graph attention network to learn the inter-connections among texts and frames. These connections incorporate both text-frame interactions and frame contextual information, which are subsequently used to aggregate the text candidates into the final enriched text embedding. To further enhance fine-grained text-frame relationship learning in FRL, EAM employs energy-based model to accurately model the energy of detailed text-frame interactions and thus capture the true distribution of text-video pairs, improving the overall matching performance. Moreover, we proposed replacing the conventional softmax-based contrastive loss [35] with sigmoid loss [56] for more effective cross-modal alignment and stable training. Softmax loss normalizes the logits across both dimensions of the batch similarity matrix, making it sensitive to the negative samples and batch size [8, 24, 56]. Instead, sigmoid loss treats each pair independently, so it can mitigate these drawbacks and is naturally suitable for the multi-matched scenario of TVR, where one text may be semantically similar to multiple videos or vice versa. Extensive experiments have been conducted on various benchmarks, demonstrating our effectiveness and superiority.
Our main contributions can be summarized as follows:
  • We design a fine-grained relationship learning network that considers not only text-video interaction but also internal frame-to-frame relations within videos, generating accurate enriched text embedding that incorporates frames’ contextual information;
  • To the best of our knowledge, we are the first to integrate EBM into TVR, which effectively models fine-grained text-frame interactions, enhancing the relational learning and cross-modal alignment.
  • Extensive experiments illustrate the superiority of EagleNet and the effectiveness of each component, showing improvement over state-of-the-art models across four benchmark datasets.
2 Related Work
Text-Video Retrieval. Text-Video Retrieval (TVR) needs to align text embeddings with video embeddings [18, 50, 47, 21]. Early works are mostly offline models that focus on designing a matching algorithm for pre-extracted features [55, 9, 49, 10, 58, 16, 14, 29, 26, 46, 5, 31, 13]. JPoSE [49] and DualEncoding [10] extract video-text features offline, then perform embedding alignment in a common space. To enhance representation quality, CE [31], MMT [14], and HiT [29] employ multiple unimodal experts (OCR, speech, etc.) to enrich video features. TeachText [9] adopts multiple text encoders to gather text information and obtain more credible alignments. Besides, HGR [5] and T2vlad [46] build semantic graphs or shared centers to explore global-local relations. However, these models employ additional analysis tools or experts to pre-process data, which increases the complexity and limits flexibility.
Recently, more and more researchers aim to seek end-to-end solutions [33, 17, 34, 53, 48, 27, 43, 12, 32, 7, 26, 16, 58]. ClipBERT [26] and Frozen [2] first propose efficient end-to-end pre-trained schemes. MCQ [16] builds cross-modal associations by predicting verb or noun features. Later, benefiting from the large-scale vision-language pre-trained models [37], CLIP4clip [33] uses CLIP as the backbone network, significantly improving TVR performance, and inspiring a series of works [32, 17, 58, 12, 22, 3, 40]. Centerclip [58] conducts segment-level clustering, which reduces token redundancy and computational overhead. TS2Net [32] designs the token shift module to realize the perception of local movement between frames. CLIP-VIP [53] learns video proxies to incorporate frame context. XPool [17] uses text as a condition to guide the aggregation of video tokens. UCoFiA [48] conducts hierarchical alignment to achieve coarse- and fine-grained matching. Similarly, GLSCL [57] proposes a global and local interaction module for coarse- and fine-grained semantic consistent learning. ProST [27] extracts higher-level prototypes for texts and videos and conducts spatio-temporal matching. TempMe [41] introduces a multi-granularity framework to progressively combine neighboring clips, reducing video redundancy and enhancing temporal modeling across frames. Video-ColBERT [38] utilizes spatiotemporal token interactions, learnable query and visual expansion, and sigmoid loss training.
However, most existing works focus on learning high-quality video embeddings or improving multi-modal alignment, while overlooking the limited expressiveness of textual representations. TMASS [43] extends text embeddings effectively by sampling stochastic text embeddings within a flexible range, whose radius is determined by the similarity between texts and videos. Similarly, TV-ProxyNet [51] converts the text query to a specific text proxy for each candidate video proxy through video-aware directors and text-video similarities. Nevertheless, these methods only consider text-video interactions while expanding text semantics, ignoring the internal frame-to-frame relations within a video, so the final enriched text cannot capture frame context, leading to disparities between text and video.
Graph Neural Networks. Graph Neural Network (GNN) has received significant interest in recent years due to its powerful ability in real-world applications based on graph-structured data [25, 42, 39]. The Graph Attention Network (GAT) [42] is the first to introduce the attention mechanism into GNN, and it shows great potential in modeling relationships. Furthermore, to tackle complex real-life problems related to heterogeneous graphs, such as processing knowledge graphs, Relational GNN (RGNN) [39] and many successors have been proposed. HGR [5] is the first work to employ GNN in TVR; it constructs semantic graphs for texts by grammar and extracts representations using GNN. Later, many following works are proposed to employ GNN in TVR [44, 23, 19]. ACP [13] uses knowledge graphs as prior knowledge to complete the semantic information. SHE-Net [54] introduces CLIP as the backbone and uses hierarchical text graphs to guide the aggregation of temporal frames and spatial regions. But these works are based on pre-extracted features and heavily rely on external knowledge, limiting the flexibility and increasing model complexity.
Unlike previous methods, our approach captures both the text-video interactions and rich relationships among video frames, which are essential for comprehending the video context. By leveraging these fine-grained interactions, the model is able to generate high-quality, context-aware text embeddings and achieve more accurate text-video alignment.
3 Method
Figure 2: Diagram of EagleNet. (a) Overview of the training process. (b) Fine-Grained Relationship Learning (FRL) first samples multiple text candidates, then constructs a text-frame graph to learn both text-frame and frame-frame relationships, which are then used to aggregate text candidates into a final enriched text embedding aware of video context information. (c) Energy-Aware Matching (EAM) improves the relationships learning in FRL by capturing detailed text-frame interactions and accurately models true text-video pairs distribution, thereby also enhancing the final matching performance from a fine-grained perspective.
3.1 Preliminary
We first introduce the standard pipeline of the text-video retrieval task, including: 1) obtaining representations of text and video; 2) training and inference.
3.1.1 Text and Video Representations
Text-video retrieval task aims to align the text and video in a joint space. The first step is to learn the text and video embeddings using the vision-language pre-trained model CLIP [37]. Specifically, given a text-video pair $\{t,v\}$, TVR first samples $M$ frames from $v$ uniformly, $f_{1},f_{2},\cdots,f_{M}]$, and then feeds them into CLIP text encoder $\phi_{\text{t}}(\cdot)$ and image encoder $\phi_{\text{v}}(\cdot)$, obtaining text embedding $\mathbf{t}$} and frame embeddings $\mathbf{F}$} as:
${\mathbf{t}}$ $=\phi_{\text{t}}(t)\in\mathbb{R}^{d},$ (1)
${\mathbf{F}}$ $=\phi_{\text{v}}([f_{1},f_{2},\cdots,f_{M}])=[{\mathbf{f}}_{1}\|{\mathbf{f}}_{2}\|\cdots\|{\mathbf{f}}_{M}]^{\top}\in\mathbb{R}^{M\times d},$
where $\|$” is the column concatenation operation, and $d$ is the hidden dimension of embeddings.
Then a frame fusion module is employed to obtain the global video feature ${\mathbf{v}}=\pi({\mathbf{F}})\in\mathbb{R}^{d}$, to reduce the temporal redundancy in video and align $\mathbf{t}$} with $\mathbf{v}$}. A simple solution is average pooling. But it treats all the frames equally, overlooking the important frames and retaining noise information from irrelevant frames. We follow XPool [17] (Appendix A) to use $\mathbf{t}$} as condition, such that the fusion module can attach more attention to the text-related frames:
${\mathbf{v}}=\pi({\mathbf{F}}|{\mathbf{t}})\in\mathbb{R}^{d},$ (2)
which uses cross-attention to fuse frames guided by text.
3.1.2 Training and Inference
In training stage, for each training batch $\{(t_{i},v_{i})\}_{i=1}^{B}$ with batch size $B$, a symmetric cross-entropy loss $\mathcal{L}_{\text{ce}}$ is applied to minimize the distance of relevant text-video pairs while maximizing that between irrelevant pairs, similar to contrastive learning [35]. We consider both text-to-video retrieval ($\mathcal{L}_{\text{ce}}^{t2v}$) and video-to-text retrieval ($\mathcal{L}_{\text{ce}}^{v2t}$):
$\mathcal{L}_{\text{ce}}({\mathbf{t}},{\mathbf{v}})$ $=\frac{1}{2}\Big(\mathcal{L}_{\text{ce}}^{t2v}({\mathbf{t}},{\mathbf{v}})+\mathcal{L}_{\text{ce}}^{v2t}({\mathbf{t}},{\mathbf{v}})\Big),$ (3)
$\mathcal{L}_{\text{ce}}^{t2v}({\mathbf{t}},{\mathbf{v}})$ $=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{e^{\operatorname{s}({\mathbf{t}}_{i},{\mathbf{v}}_{i})\cdot\lambda}}{\sum_{j=1}^{B}e^{\operatorname{s}({\mathbf{t}}_{i},{\mathbf{v}}_{j})\cdot\lambda}},$
$\mathcal{L}_{\text{ce}}^{v2t}({\mathbf{t}},{\mathbf{v}})$ $=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{e^{\operatorname{s}({\mathbf{t}}_{i},{\mathbf{v}}_{i})\cdot\lambda}}{\sum_{j=1}^{B}e^{\operatorname{s}({{\mathbf{t}}_{j}},{\mathbf{v}}_{i})\cdot\lambda}},$
where $\operatorname{s}(\cdot,\cdot)$ is similarity function (e.g., cosine similarity) and $\lambda$ is a learnable factor.
In inference, given $n_{t}$ texts $\{t_{i}\}_{i=1}^{n_{t}}$ and $n_{v}$ video clips $\{v_{j}\}_{j=1}^{n_{v}}$, for a query text $t_{\text{query}}$, the matched video clip $v^{*}$ is the one that has the highest matching score and vice versa:
$v^{*}=\operatorname*{{\arg\max}}_{j}\operatorname{s}({\mathbf{t}}_{\text{query}},{\mathbf{v}}_{j}),\ \ \ t^{*}=\operatorname*{{\arg\max}}_{i}\operatorname{s}({\mathbf{t}}_{i},{\mathbf{v}}_{\text{query}}).$
3.2 Our Proposed Model: EagleNet
In this section we introduce EagleNet (Fig. 2), which comprises two key components: 1) Fine-Grained Relationship Learning (FRL) and 2) Energy-Aware Matching (EAM). First, FRL generates context-aware enriched text by sampling text candidates and extracting fine-grained relationships among texts and frames (Sec. 3.2.1). Then, EAM enhances FRL in learning text-frame interactions and improves the final matching from a fine-grained view (Sec. 3.2.2). Finally, the sigmoid function is employed in the final loss for more effective alignment and stable training (Sec. 3.2.3).
3.2.1 Fine-Grained Relationship Learning
Existing models only consider text-frame interactions when expanding text semantics. However, the internal frame-frame relations are also important to capture both global and temporal semantics. Specifically, we first sample multiple text candidates, and the texts and frames are treated as nodes to form a text-frame graph, which includes both text-frame and frame-frame relations. Then the relations are extracted by relational graph learning and used to fuse text candidates, generating text embedding that reflects not only the semantics of individual frames but also their contextual dependencies.
Text-Frame Graph Construction. Given text embedding $\mathbf{t}$}} and frame embeddings ${\mathbf{F}}=[{\mathbf{f}}_{1}\|{\mathbf{f}}_{2}\|\cdots\|{\mathbf{f}}_{M}]^{\top}$, we first sample $S$ stochastic text candidates $\{{{\mathbf{t}}_{i}^{\text{sto}}}\}_{i=1}^{S}$ using stochastic text modeling [43] (Appendix B), then concatenate texts and frames to get the node matrix as:
$\mathbf{X}$ $=\left[{{\mathbf{t}}}\|{{\mathbf{t}}_{1}^{\text{sto}}}\|{{\mathbf{t}}_{2}^{\text{sto}}}\|\cdots\|{{\mathbf{t}}_{S}^{\text{sto}}}\|{\mathbf{f}}_{1}^{\prime}\|{\mathbf{f}}_{2}^{\prime}\|\cdots\|{\mathbf{f}}_{M}^{\prime}\right]^{\top}\in\mathbb{R}^{n\times d},$ (4)
${\mathbf{f}}_{j}^{\prime}$ $={\mathbf{f}}_{j}+\operatorname{PE}(j),$
where $n=1+S+M$ is the number of nodes, and $\operatorname{PE}(\cdot)$ is learnable temporal positional embedding on frames. Thus, there are three types of edges (relations): text-text, frame-frame, and text-frame, denoted by $r\in\{tt,ff,tf\}$.
Relationship Learning. To extract useful information from the complex relations among texts and frames, we propose a simple yet effective version of the Relational Graph Attention Network (RGAT) with $L$ layers. The layer index is omitted here for brevity. In each layer, for any node pair $(i,j)$ linked by relation $r$, the edge weight is given by
$e_{ij}^{r,h}$ $=\psi^{r}\left(\left[\mathbf{W}^{r,h}\mathbf{h}_{i}\|\mathbf{W}^{r,h}\mathbf{h}_{j}\right]\right)\in\mathbb{R},$ (5)
where $h=1,2,\cdots,H$ is the indices of attention head, $\mathbf{h}_{i}$ is the feature of node $i$ (in the first layer $\mathbf{h}_{i}$ is the $i$-th row of $\mathbf{X}$), $\mathbf{W}^{r,h}$ is learnable matrix, and $\psi^{r}(\cdot):\mathbb{R}^{2d}\mapsto\mathbb{R}$ is projector to obtain edge weights. Then, the attention score is obtained as
$\alpha_{ij}^{r,h}$ $=\frac{\exp\left(\operatorname{LeakyReLU}(e_{ij}^{r,h})\right)}{\sum_{k\in\mathcal{N}_{i}^{r}}\exp\left(\operatorname{LeakyReLU}(e_{ik}^{r,h})\right)},$ (6)
where $\mathcal{N}_{i}^{r}$ is the neighbor set of node $i$ linked by relation $r$. Finally, the output node representation of this layer $\mathbf{h}_{i}^{\prime}$ is
$\mathbf{h}_{i}^{\prime}=\operatorname{ReLU}\Big(\mathbf{W}_{\text{out}}\mathbf{h}_{i}+\Big\|_{h=1}^{H}\sum_{r}\sum_{j\in\mathcal{N}_{i}^{r}}\alpha_{ij}^{r,h}\mathbf{W}^{r,h}\mathbf{h}_{j}\Big),$ (7)
where $\mathbf{W}_{\text{out}}\in\mathbb{R}^{Hd\times d}$ is learnable weights for residual connection. Then $\mathbf{h}_{i}^{\prime}$ is input of the next layer. Matrix form and more details can be found in Appendix C.
Note that here $\mathbf{h}_{i}^{\prime}$ only serves as an intermediate. The target of FRL is text-frame relations $e_{ij}^{tf}$, which are regarded as the importance of text embeddings for the frames. These weights are then used to aggregate text embeddings into a final enriched text. Specifically, the text-frame edge weights from the final layer are averaged across attention heads to get the final text-frame weights, i.e.,
$e_{ij}^{tf}=\frac{1}{H}\sum_{h=1}^{H}e_{ij}^{tf,h},\ \ \begin{aligned} &i\in[1,1+S],\\ &j\in[1+S+1,1+S+M],\end{aligned}$ (8)
then we average and normalize the text-frame weights to obtain the attention score of each text node:
$w_{i}=\frac{\exp(e_{i}^{tf})}{\sum_{k=1}^{1+S}\exp(e_{k}^{tf})},\ \ \ e_{i}^{tf}=\frac{1}{M}\sum_{j=1+S+1}^{1+S+M}e_{ij}^{tf},$ (9)
so the final enriched text embedding is
${\mathbf{t}}^{\text{gen}}=\sum_{i=1}^{1+S}w_{i}\mathbf{X}_{i}=w_{1}{{\mathbf{t}}}+\sum_{i=1}^{S}w_{i+1}{{\mathbf{t}}_{i}^{\text{sto}}}\in\mathbb{R}^{d}.$ (10)
${\mathbf{t}}^{\text{gen}}$ is used as the condition for frame fusion (Eq. 2) to obtain video embedding $\mathbf{v}$}, then $\mathcal{L}_{\text{ce}}({\mathbf{t}}^{\text{gen}},{\mathbf{v}})$ is calculated (Eq. 3). Following [43] (Appendix B), we further obtain a support text ${\mathbf{t}}^{\text{sup}}$, and calculate the support loss $\mathcal{L}_{\text{ce}}({\mathbf{t}}^{\text{sup}},{\mathbf{v}})$. ${\mathbf{t}}^{\text{sup}}$ serves as the boundary of the stochastic modeling range, within which ${\mathbf{t}}_{i}^{\text{sto}}$} is sampled. So the matching loss is
$\mathcal{L}_{\text{match}}=\mathcal{L}_{\text{ce}}({\mathbf{t}}^{\text{gen}},{\mathbf{v}})+\lambda_{\text{sup}}\mathcal{L}_{\text{ce}}({\mathbf{t}}^{\text{sup}},{\mathbf{v}}),$ (11)
where $\lambda_{\text{sup}}$ is a coefficient.
3.2.2 Energy-Aware Matching
Method Text-to-Video Video-to-Text Rsum↑
R@1↑ R@5↑ R@10↑ MdR↓ MnR↓ R@1↑ R@5↑ R@10↑ MdR↓ MnR↓
Not End-to-End Method
CE [31] BMVC19 20.6 50.3 64.0 5.3 - 20.9 48.8 62.4 6.0 - 267.0
MMT [14] ECCV20 26.6 57.1 69.6 4.0 24.0 27.0 57.5 69.7 3.7 21.3 307.5
T2vlad [46] CVPR21 29.5 59.0 70.1 4.0 - 31.8 60.0 71.1 3.0 - 321.5
HiT [29] ICCV21 27.7 59.2 72.0 2.9 - 28.8 60.3 72.3 3.0 - 320.3
TeachText [9] ICCV21 29.6 61.6 74.2 3.0 - 32.1 62.7 75.0 3.0 - 335.2
End-to-End Method
non-CLIP ClipBERT [26] CVPR21 22.0 46.8 59.9 6.0 - - - - - - -
SupportSet [36] ICLR21 30.1 58.5 69.3 3.0 - 28.5 58.6 71.6 3.0 - 316.6
Frozen [2] ICCV21 31.0 59.5 70.5 3.0 - - - - - - -
BridgeFormer [16] CVPR22 37.6 64.8 75.1 3.0 - - - - - - -
TVMM [28] NeurIPS22 36.2 64.2 75.7 3.0 - 34.8 63.8 73.7 3.0 - 348.4
ViT-B/32 CLIP4Clip [33] NeurCom22 43.2 71.4 80.6 2.0 16.5 42.5 70.9 81.0 2.0 12.2 389.6
XPool†∗ [17] CVPR22 45.8 73.8 82.0 2.0 14.0 45.7 73.1 82.3 2.0 9.4 402.7
TS2Net [32] ECCV22 45.4 72.9 82.9 2.0 14.2 44.8 72.6 81.1 2.0 10.3 399.7
CLIP-VIP†∗ [53] ICLR23 46.0 72.6 82.2 2.0 14.4 43.6 71.5 82.3 2.0 10.8 398.2
ProST [27] ICCV23 46.3 73.6 82.4 2.0 12.9 45.0 72.9 82.2 2.0 8.9 402.4
TMASS†∗ [43] CVPR24 43.4 73.2 82.1 2.0 14.3 45.0 73.2 82.7 2.0 9.5 399.6
RAP [3] ACL24 44.8 71.4 81.5 - 14.4 44.0 71.9 82.4 - 10.1 396.0
TFVL [40] TNNLS25 45.2 70.4 80.1 - 15.2 45.5 71.5 81.0 - 11.0 393.7
SHE-Net [54] TCSVT25 45.3 74.9 84.2 2.0 12.5 44.8 74.8 83.7 2.0 9.1 407.7
GLSCL [57] TIP25 47.2 73.5 82.7 2.0 13.1 45.6 73.5 82.7 2.0 9.8 405.2
TempMe [41] ICLR25 43.3 71.2 81.2 2.0 15.3 45.0 71.5 80.8 2.0 11.0 393.0
TV-ProxyNet†∗ [51] AAAI25 46.7 73.0 81.8 2.0 14.5 45.4 71.1 82.4 2.0 11.0 400.4
Video-ColBERT†∗ [38] CVPR25 45.9 71.6 82.7 2.0 13.7 42.1 70.4 81.1 2.0 10.8 393.8
EagleNet (Ours) 47.2 74.8 84.0 2.0 14.4 45.7 73.7 83.4 2.0 9.4 408.8
ViT-B/16 CLIP4Clip [33] NeurCom22 45.2 72.2 81.4 2.0 15.0 42.9 70.9 80.6 2.0 11.6 393.2
XPool†∗ [17] CVPR22 49.2 73.9 82.6 2.0 13.6 48.0 74.9 82.9 2.0 9.7 411.5
TS2Net [32] ECCV22 48.8 75.0 85.0 2.0 13.0 46.4 75.4 85.3 2.0 8.6 415.9
CLIP-VIP†∗ [53] ICLR23 48.5 75.5 83.9 2.0 12.5 47.6 74.8 85.0 2.0 9.6 415.3
UCoFiA [48] ICCV23 48.6 75.5 83.8 2.0 11.7 49.0 75.7 84.2 2.0 10.0 416.8
ProST [27] ICCV23 48.0 74.7 84.0 2.0 12.7 48.3 75.2 84.5 2.0 8.7 414.7
TMASS†∗ [43] CVPR24 48.5 74.1 83.4 2.0 14.4 46.7 75.9 85.6 2.0 9.1 414.2
RAP [3] ACL24 46.5 73.9 82.0 - 12.1 45.3 76.4 84.8 - 9.1 408.9
TFVL [40] TNNLS25 48.7 73.5 83.0 - 13.4 47.3 75.3 83.7 - 9.5 411.5
GLSCL [57] TIP25 49.9 76.3 84.1 2.0 11.9 48.3 75.8 84.6 2.0 8.1 419.0
TempMe [41] ICLR25 49.9 75.3 84.8 2.0 12.1 47.5 75.2 85.1 2.0 8.8 417.8
TV-ProxyNet†∗ [51] AAAI25 49.5 74.9 83.4 2.0 12.6 47.9 74.8 84.1 2.0 9.4 414.6
Video-ColBERT†∗ [38] CVPR25 50.0 76.3 84.3 1.5 13.1 47.9 74.4 84.9 2.0 9.0 417.8
EagleNet (Ours) 51.0 76.2 85.6 1.0 12.8 49.2 77.1 86.6 2.0 8.2 425.7
Table 1: Performance on MSRVTT, boldface is the best results and underline is the runner-ups. Note that we have resolved the data-leakage problem in the original TMASS codes, and re-implemented XPool, CLIP-VIP, TMASS, TV-ProxyNet, and Video-ColBERT using CLIP4Clip codebase to align with other methods. denotes the reproduced results, all of which are fine-tuned on ViT-B/32 or ViT-B/16.
Since $\mathcal{L}_{\text{match}}$ only aligns text and video at global level, we propose EAM to assist FRL in learning fine-grained text-frame relations. EAM employs energy-based model to effectively capture data distributions [11, 30, 6], thereby enhancing the matching from a fine-grained perspective.
To be specific, given text embedding $\mathbf{t}$}} and frame embeddings ${\mathbf{F}}=[{\mathbf{f}}_{1}\|\cdots\|{\mathbf{f}}_{M}]^{\top}$, EAM models the real text-video pair distribution $p({{\mathbf{t}}},{\mathbf{F}})$ using Boltzmann distribution $p_{\theta}({{\mathbf{t}}},{\mathbf{F}})=\frac{\exp{(-E_{\theta}({{\mathbf{t}}},{\mathbf{F}}))}}{Z_{\theta}}$, where $Z_{\theta}$ is the normalizing constant, $\theta$ is learnable parameter, and $E_{\theta}({{\mathbf{t}}},{\mathbf{F}})\in\mathbb{R}$ denotes the text-video energy. We first calculate the energy for each text-frame pair and then average them to obtain the text-video energy $E_{\theta}({{\mathbf{t}}},{\mathbf{F}})=\frac{1}{M}\sum_{i}^{M}E_{\theta}({{\mathbf{t}}},{\mathbf{f}}_{i})$. Such that EAM can fully utilize the information in fine-grained text-frame interactions. Simultaneously, $E_{\theta}$ assigns low energy scores to the true text-video pairs and high energy to mismatched ones, further assisting the model in achieving more precise text-video alignment.
Additionally, besides averaging text-frame energies (Avgpool), other pooling approaches, e.g., Maxpool, are also options. This will be discussed in Sec. 4.3.3.
Energy Functions. Since the energy score indicates the relevance of a text-video pair, we can adopt a similarity function as $E_{\theta}$, such as negative cosine similarity (CosSim):
$E({{\mathbf{t}}},{\mathbf{F}})=\frac{1}{M}\sum_{i=1}^{M}E({{\mathbf{t}}},{\mathbf{f}}_{i})=-\frac{1}{M}\sum_{i=1}^{M}\frac{{{\mathbf{t}}}^{\top}{\mathbf{f}}_{i}}{\|{{\mathbf{t}}}\|\|{\mathbf{f}}_{i}\|}.$ (12)
Or we can use the bilinear scoring function with learnable parameters $\mathbf{W}$ (Bilinear):
$E_{\theta}({{\mathbf{t}}},{\mathbf{F}})=-\frac{1}{M}\sum_{i=1}^{M}\frac{{{{\mathbf{t}}}}^{\top}\mathbf{W}{\mathbf{f}}_{i}}{\|{{\mathbf{t}}}\|\|{\mathbf{f}}_{i}\|}.$ (13)
For greater flexibility, MLP is also a strong choice (MLP):
$E_{\theta}({{\mathbf{t}}},{\mathbf{F}})=\frac{1}{M}\sum_{i=1}^{M}\operatorname{MLP}([{{\mathbf{t}}}\|{\mathbf{f}}_{i}]).$ (14)
Method DiDeMo MSVD VATEX Rsum↑
R@1↑ R@5↑ R@10↑ R@1↑ R@5↑ R@10↑ R@1↑ R@5↑ R@10↑
CE [31] 16.1 41.1 82.7 19.8 49.0 - - - - -
HGR [5] - - - - - - 35.1 73.5 83.5 -
ClipBERT [26] 20.4 48.0 60.8 - - - - - - -
SupportSet [36] - - - 28.4 60.0 - 45.9 82.4 - -
Frozen [2] 31.0 59.8 72.4 33.7 64.7 - - - - -
BridgeFormer [16] 37.0 62.2 73.9 - - - - - - -
TVMM [28] 36.5 64.9 75.4 - - - - - - -
ViT-B/32 CLIP4Clip [33] 40.6 67.7 77.8 46.2 75.6 84.6 56.6 87.5 93.6 630.2
XPool†∗ [17] 43.5 69.5 79.0 47.1 76.5 85.2 57.5 87.8 93.7 639.8
TS2Net [32] 39.5 69.7 80.9 44.7 75.5 84.9 57.7 88.7 94.2 635.8
CLIP-VIP†∗ [53] 44.7 72.3 81.3 46.5 76.2 85.2 59.3 89.0 94.3 648.8
ProST [27] 45.8 74.3 84.5 44.3 75.2 84.8 57.3 88.1 94.2 648.5
TMASS†∗ [43] 43.4 70.6 81.0 41.8 71.9 81.4 56.1 88.1 93.8 628.1
RAP [3] 42.6 70.4 79.6 44.9 73.7 83.1 - - - -
SHE-Net [54] 45.6 75.6 83.6 47.6 76.8 85.5 - - - -
GLSCL [57] 45.5 75.4 84.3 - - - - - - -
TempMe [41] 47.2 73.7 81.3 - - - - - - -
TV-ProxyNet†∗ [51] 46.8 72.1 81.3 46.6 76.0 85.0 58.9 89.1 94.4 650.2
Video-ColBERT†∗ [38] 40.2 68.6 79.3 43.9 73.9 83.5 57.1 87.4 93.1 627.0
EagleNet (Ours) 46.8 74.0 83.5 47.9 77.7 85.9 59.0 89.2 94.1 658.1
ViT-B/16 CLIP4Clip [33] 45.1 71.5 81.4 47.6 78.0 86.4 60.9 90.6 95.4 656.9
XPool†∗ [17] 44.1 73.5 83.4 49.5 79.6 87.9 62.0 91.0 95.6 666.6
TS2Net [32] 43.1 72.3 81.5 47.0 78.0 86.6 62.1 91.2 95.8 657.6
CLIP-VIP†∗ [53] 47.2 75.5 84.4 49.2 79.1 87.5 63.8 91.6 96.0 674.3
ProST [27] 47.0 75.5 84.2 48.1 78.6 87.4 62.0 91.3 95.8 669.9
TMASS†∗ [43] 42.1 73.1 82.3 44.9 75.2 84.2 60.8 90.5 95.3 648.4
GLSCL [57] 49.1 76.5 85.7 - - - - - - -
TempMe [41] 50.2 76.4 85.0 - - - - - - -
TV-ProxyNet†∗ [51] 47.9 75.2 84.5 49.7 79.7 88.0 64.0 91.7 95.9 676.6
Video-ColBERT†∗ [38] 44.5 71.7 82.1 47.7 78.3 86.6 61.1 90.2 95.3 657.5
EagleNet (Ours) 51.5 78.7 86.8 50.9 80.7 88.3 63.6 91.5 95.7 687.7
Table 2: Text-to-video results on DiDeMo, MSVD, and VATEX. Note that we have resolved the data-leakage problem in the original TMASS codes, and re-implemented XPool, CLIP-VIP, TMASS, TV-ProxyNet, and Video-ColBERT using CLIP4Clip codebase to align with other methods. denotes the reproduced results, all of which are fine-tuned on ViT-B/32 or ViT-B/16.
Training EAM. In practice, EAM is trained by $\mathcal{L}_{\text{eam}}=-\mathbb{E}_{\{{{\mathbf{t}}},{\mathbf{F}}\}\sim p({{\mathbf{t}}},{\mathbf{F}})}[\log p_{\theta}({{\mathbf{t}}},{\mathbf{F}})]$, this negative log-likelihood loss can be reformulated as:
$\mathcal{L}_{\text{eam}}\approx\frac{1}{B}\sum_{i=1}^{B}E_{\theta}({\mathbf{t}}_{i},{\mathbf{F}}_{i})-\frac{1}{B}\sum_{j=1}^{B}E_{\theta}(\tilde{{{\mathbf{t}}}}_{j},\tilde{{\mathbf{F}}}_{j}),$ (15)
[…]
where $\tilde{{{\mathbf{t}}}}_{j}^{(0)}\in\mathbb{R}^{d}$ and $\tilde{{\mathbf{F}}}_{j}^{(0)}\in\mathbb{R}^{M\times d}$ are given initial samples, $\eta$ is the step size, and $\bm{\bm{\epsilon}}^{(k)}\sim\mathcal{N}\left(0,\sigma^{2}\right)$ with a given noise $\sigma^{2}$. We stop the gradients of $\tilde{{{\mathbf{t}}}}$ and $\tilde{{\mathbf{F}}}$ after sampling. More details of the derivation and sampling algorithm can be found in Appendix D. EAM is only involved in the training process, and thus will not increase inference cost.
We incorporate $\mathcal{L}_{\text{eam}}$ to define the total objective:
$\mathcal{L}_{\text{total}}=$ $\mathcal{L}_{\text{match}}+\lambda_{\text{eam}}\mathcal{L}_{\text{eam}}$ (17)
$=$ $\mathcal{L}_{\text{ce}}({\mathbf{t}}^{\text{gen}},{\mathbf{v}})+\lambda_{\text{sup}}\mathcal{L}_{\text{ce}}({\mathbf{t}}^{\text{sup}},{\mathbf{v}})+\lambda_{\text{eam}}\mathcal{L}_{\text{eam}},$
where $\lambda_{\text{eam}}$ is a coefficient.
3.2.3 Sigmoid Loss
The softmax-based contrastive loss ($\mathcal{L}_{\text{ce}}$) is sensitive to negative samples and batch size due to the batch normalization [8, 24, 56]. Instead, the sigmoid loss [56] models each text-video pair independently, enabling more effective alignment and stable optimization, so is naturally suitable for the multi-matched scenario of TVR, where one text may be semantically similar to multiple videos or vice versa.
$\mathcal{L}_{\text{sig}}({\mathbf{t}},{\mathbf{v}})=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{B}\log\frac{1}{1+e^{\mathbb{I}_{ij}\left(\tau\cdot\operatorname{s}({\mathbf{t}}_{i},{\mathbf{v}}_{j})+b\right)}}$ (18)
where $\mathbb{I}_{ij}$ is a indicator for positive pairs (+1+1) and negative pairs (1-1), $\tau=e^{\tau_{p}}$, $\tau_{p}$ and $b$ are learnable and initialized as $\tau_{p}=4.77$ and b=12.93b=-12.93, according to the pre-training in [56]. Finally, we replace $\mathcal{L}_{\text{ce}}$ with $\mathcal{L}_{\text{sig}}$ in Eq. 17:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{sig}}({\mathbf{t}}^{\text{gen}},{\mathbf{v}})+\lambda_{\text{sup}}\mathcal{L}_{\text{sig}}({\mathbf{t}}^{\text{sup}},{\mathbf{v}})+\lambda_{\text{eam}}\mathcal{L}_{\text{eam}}.$ (19)
4 Experiment
4.1 Experimental Settings
Datasets and Metrics. We use four benchmark datasets: MSRVTT [52], DiDeMo [1], MSVD [4], and VATEX [45]. These datasets contain sufficient video clips, each of which has multiple captions. We use the standard TVR metrics [33]: Recall (%) at rank 1/5/10 (R@1/5/10), Median Rank (MdR), Mean Rank (MnR), and Rsum (R@1++R@5++R@10); and also the total test time (sec.). More details can be found in Appendix E.
Implementation. We initialized our model by CLIP pre-trained models (ViT-B/32 and ViT-B/16) [37], so hidden dimension $d=512$. We uniformly sample $M=12$ frames per video ($64$ for DiDeMo) and the max length of text is $32$ ($64$ for DiDeMo). In FRL, we sample $S=20$ text candidates. For MCMC sampling in EAM, step number $K=20$, step size $\eta=1$, and noise variance $\sigma^{2}=0.005$. For loss coefficient, we set $\lambda_{\text{sup}}=0.8$ following [43] and $\lambda_{\text{eam}}=1.0$ by default. We train on two A100 GPUs for $5$ epochs with batch size $B=64$, weight decay $0.2$, dropout $0.3$, learning rate $10^{-7}$ for CLIP modules and $10^{-4}$ for non-CLIP modules; and we evaluate on one A100 with batch size $8$. More details can be found in Appendix E.
4.2 Comparison to State-of-the-art Models
We compare our proposed EagleNet with existing state-of-the-art (SOTA) models [33, 17, 32, 53, 48, 27, 43, 57, 41, 51, 38]. We show the performance in Tabs. 1 and 2, and the test time in Fig. 1. See Appendix F for further analysis and complexity details. Note that we have addressed the data-leakage problem in the original TMASS codes, and re-implemented XPool, CLIP-VIP, TMASS, TV-ProxyNet, and Video-ColBERT using CLIP4Clip codebase to align with other models for fair comparisons.
Overall, EagleNet achieves the best text-to-video R@1 performance across datasets, outperforming the second-best by 2.00% on MSRVTT, 2.59% on DiDeMo, and 2.41% on MSVD, respectively. And our model consistently reaches the highest Rsum results, demonstrating the superiority and universality of our proposed model.
Figure 3: Visualization of EagleNet, TMASS, and TV-ProxyNet.
Components MSRVTT DiDeMo Avg
FRL EAM $\mathcal{L}_{\text{sig}}$ R@1↑ R@5↑ R@10↑ R@1↑ R@5↑ R@10↑ R@1↑ R@5↑ R@10↑
48.5 74.1 83.4 42.1 73.1 82.3 45.3 73.6 82.9
48.8 75.6 83.9 47.9 75.2 84.1 48.4 75.4 84.0
49.0 75.3 84.0 43.4 73.6 82.7 46.2 74.5 83.4
50.5 76.5 84.2 49.2 76.6 85.3 49.9 76.6 84.8
47.8 75.9 84.4 43.9 75.6 83.6 45.9 75.8 84.0
49.9 77.3 84.3 49.1 77.5 85.0 49.5 77.4 84.7
48.6 75.3 83.7 43.8 75.6 83.8 46.2 75.5 83.8
EagleNet 51.0 76.2 85.6 51.5 78.7 86.8 51.3 77.5 86.2
Table 3: Text-to-video results on different components.
Energy MSRVTT DiDeMo
Function R@1↑ R@5↑ R@10↑ R@1↑ R@5↑ R@10↑
CosSim 50.0 76.9 85.5 50.2 78.3 86.3
Bilinear 51.0 76.2 85.6 51.3 78.3 86.6
MLP 50.3 76.6 85.1 51.5 78.7 86.8
Table 4: Text-to-video on different energy functions.
Pooling MSRVTT DiDeMo
Form R@1↑ R@5↑ R@10↑ R@1↑ R@5↑ R@10↑
w/o EAM 49.9 77.3 84.3 49.1 77.5 85.0
$E_{\theta}({\mathbf{t}},{\mathbf{v}})$ 48.8 76.3 84.6 47.8 77.1 85.2
Maxpool 50.4 76.5 85.5 51.3 79.0 86.6
Minpool 50.5 77.2 84.9 50.9 78.5 86.5
Avgpool 51.0 76.2 85.6 51.5 78.7 86.8
Table 5: Text-to-video on different pooling approach for $E_{\theta}$.
Both TMASS [43] and TV-ProxyNet [51] focus solely on the interactions between text and frames/video, so they fail to exploit temporal information within the video, resulting in suboptimal performance (MSRVTT text-to-video R@1 48.5% for TMASS and 49.5% for TV-ProxyNet). From Fig. 3, they might retrieve incorrect videos with similar semantics but different contexts. In contrast, EagleNet also leverages the internal frame-to-frame interactions within the video when expanding text semantics. Modeling such frame-level relationships enables the enriched text embedding to capture both the global video semantics and the temporal contextual information, leading to SOTA performance across benchmarks. Visualization of not considering frame context (without FRL) is presented in Appendix F.5.2. Moreover, TMASS, based on XPool [17], designs a stochastic text modeling strategy to sample text embeddings with richer semantics. However, after addressing the data-leakage issue, its text-to-video R@1 result (48.5%) is worse than that of XPool (49.2%) on MSRVTT, underperforms the runner-up by 16.14% on DiDeMo, and achieves unsatisfactory performance on MSVD (44.9%) and VATEX (60.8%). These results suggest that stochastically sampling a single text embedding per text-video pair may introduce undesired noise, degrading the semantic quality of the final text representation. EagleNet instead extracts frame-relevant information from multiple text candidates, producing an accurate text embedding. We provide detailed visualization comparisons of different models in Appendix F.5.1.
4.3 Ablation Study
4.3.1 The Effect of Different Components
From Tab. 3, we first analyze the effect of different components (FRL and EAM). The base model (Row 1, R@1 48.5% on MSRVTT and 42.1% on DiDeMo) only uses the interactions between text and frames, yielding worse results than the baseline, XPool (49.2% on MSRVTT and 44.1% on DiDeMo). In contrast, FRL (Row 2) incorporates frame-frame relationships and thus captures the frame contextual information, improving the final performance (R@1 48.8% on MSRVTT and 47.9% on DiDeMo). Moreover, as mentioned in Sec. 4.2, the stochastic text modeling of TMASS (Row 1) may bring undesired noise, thereby also lowering the performance compared to its baseline, XPool. FRL instead extracts frame-related information from multiple text candidates, resulting in an accurate enriched text embedding. Meanwhile, EAM models the distribution of real text-video pairs from a fine-grained perspective, enhancing the ability of the model to consider fine-grained matching relevance and consequently improving the performance (Row 3, 49.0% on average). Combining them, EAM can enhance FRL’s performance (Row 4, 49.9% on average) in learning relationships by utilizing the fine-grained information in text-frame interactions.
Moreover, we observe an overall improvement when using $\mathcal{L}_{\text{sig}}$ over $\mathcal{L}_{\text{ce}}$ (Row 5 to 8). Since sigmoid loss models each text-video pair independently, it is unaffected by the batch size and negative samples, enabling more effective alignment and stable optimization. These features make it more suitable for TVR scenarios.
Detailed visualization comparison of different components can be found in Appendix F.5.2.
4.3.2 The Selection of Energy Function
We compare the performance of different energy functions described in Sec. 3.2.2 (Tab. 4). On the dataset MSRVTT, the Bilinear shows the best R@1 results (51.0%). On DiDeMo, the best result (51.5%) is given by MLP. This indicates that we need to select a suitable one for different applications.
Although CosSim is learning-free and shows comparable performance, such a non-parametric energy function limits the capability of dealing with complicated scenarios. The Bilinear can be regarded as CosSim with learnable parameters and is more flexible. MLP adaptively learn the connection between text and video and is no longer limited to a cosine similarity form, so it can discover more implicit information in some cases but has more leanable parameters. Overall, we choose Bilinear as our final implementation.
4.3.3 The Pooling Approach for Text-Frame Energy
Here we discuss the pooling approach for energy function, Avgpool ($\frac{1}{M}\sum_{i=1}^{M}E_{\theta}({\mathbf{t}},{\mathbf{f}}_{i})$), Maxpool ($\max_{i}E_{\theta}({\mathbf{t}},{\mathbf{f}}_{i})$), and Minpool ($\min_{i}E_{\theta}({\mathbf{t}},{\mathbf{f}}_{i})$). Since the energy of a matched pair is low, Minpool selects the frame that is most relevant to the given text. On the contrary, Maxpool selects the most irrelevant frame, which can be seen as a lower bound of text-video relevance. Besides, we can also use the global video embedding to calculate the text-video energy ($E_{\theta}({\mathbf{t}},{\mathbf{v}})$).
From Tab. 5, Avgpool shows the best R@1 performance on both MSRVTT (51.0%) and DiDeMo (51.5%), suggesting that considering all the text-frame interactions brings the most benefit to the retrieval. Regarding the other two approaches, selecting the most relevant frame yields better results on MSRVTT (Minpool, 50.5%); but improving the lower bound to enhance the overall relevance is more suitable for DiDeMo (Maxpool, 51.3%).
Besides, fusing frames before calculating energy shows lower performance (48.8% and 47.8%) than not using EAM (49.9% and 49.1%). Therefore, learning from fine-grained text-frame interactions can better improve the model in identifying the actual text-video pair. The fused video embedding provides only a global view, missing fine-grained details.
5 Conclusions
In this paper, we present EagleNet, a novel text-video retrieval framework that generates accurate and context-aware enriched text embeddings by exploring fine-grained relationships among text and video frames. We further employ an energy-based model to capture detailed interactions, which facilitates learning of relationships and enhances matching performance from a fine-grained perspective. Moreover, we instead use sigmoid loss for more effective alignment and stable training. Extensive experiments across four datasets demonstrate the superior performance of EagleNet compared with other state-of-the-art methods.
Acknowledgements
This work was supported in part by the Shenzhen Science and Technology Program under Grant No. KQTD20221101093559018, Grant No. CJGJZD20240729141505007, and Grant No. SYSRD20250529113401002, by the National Natural Science Foundation of China (NSFC) under Grant No. 62441619 and Grant No. 62411540034.
Supplementary Material
Appendix A Fusion Module for Frames
After obtaining text embedding ${{\mathbf{t}}}\in\mathbb{R}^{d}$ and frame embeddings ${\mathbf{F}}=[{\mathbf{f}}_{1}\|{\mathbf{f}}_{1}\|\cdots\|{\mathbf{f}}_{M}]^{\top}\in\mathbb{R}^{M\times d}$ for a text-video pair, we follow XPool [17] to use $\mathbf{t}$}} as condition, such that the fusion module $\pi({\mathbf{F}}|{{\mathbf{t}}})$ can attach more attention to the text-related frames.
Specifically, we employ cross-attention, where the text is the query and the frames are the key and value:
$\mathbf{Q}_{t}$ $={{\mathbf{t}}}^{\top}\mathbf{W}_{Q}\in\mathbb{R}^{1\times d_{p}},$
$\mathbf{K}_{F}$ $={\mathbf{F}}\mathbf{W}_{K}\in\mathbb{R}^{M\times d_{p}},$
$\mathbf{V}_{F}$ $={\mathbf{F}}\mathbf{W}_{V}\in\mathbb{R}^{M\times d_{p}},$
where $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$ are learnable matrices, $d_{p}$ is the dimension of the cross-attention module. Then the global video presentation $\mathbf{v}$}} can be obtained as:
$\mathbf{z}$ $=\operatorname{LN}(\operatorname{Attention}(\mathbf{Q}_{t},\mathbf{K}_{F},\mathbf{V}_{F})\mathbf{W}_{O})\in\mathbb{R}^{1\times d},$
${{\mathbf{v}}}$ $=\operatorname{LN}(\operatorname{FC}(\mathbf{z})+\mathbf{z})^{\top}\in\mathbb{R}^{d},$
where $\operatorname{LN}(\cdot)$ is the layer-norm layer, and $\operatorname{FC}(\cdot)$ is a fully-connected layer. So we have $\pi({\mathbf{F}}|{{\mathbf{t}}})={{\mathbf{v}}}$.
Appendix B Stochastic Text Embeddings Sampling
B.1 Sampling
Given text embedding ${{\mathbf{t}}}\in\mathbb{R}^{d}$ and frame embeddings ${\mathbf{F}}\in\mathbb{R}^{M\times d}$ for a text-video pair, we follow TMASS [43] to sample $S$ stochastic text embeddings $\{{{\mathbf{t}}}_{i}^{\text{sto}}\}_{i=1}^{S}$ with extended semantics.
Here we only introduce how to sample one text embedding. Specifically, we first calculate a ‘radius’ $\mathbf{r}$ of sampling:
$\mathbf{s}$ $=\operatorname{CosSim}({\mathbf{F}},{{\mathbf{t}}})\in\mathbb{R}^{M},$
$\mathbf{r}$ $=\exp(\mathbf{s}^{\top}\mathbf{W})\in\mathbb{R}^{d},$
where $\operatorname{CosSim}(\cdot,\cdot)$ is the consine similarity operation, and $\mathbf{W}\in\mathbb{R}^{M\times d}$ is a learnable matrix. Then the sampled text embedding is:
${{\mathbf{t}}}_{i}^{\text{sto}}={{\mathbf{t}}}+\mathbf{r}\odot\bm{\epsilon}_{i},$
where $\bm{\epsilon}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$. Repeat for $S=20$ times, then we can obtain $S$ stochastic text embeddings $\{{{\mathbf{t}}}_{i}^{\text{sto}}\}_{i=1}^{S}$.
B.2 Support Text
Following [43], we further obtain a support text ${\mathbf{t}}^{\text{sup}}$. ${\mathbf{t}}^{\text{sup}}$ serves as the boundary of the randomness, within whose range the stochastic text embeddings are sampled. Including ${\mathbf{t}}^{\text{sup}}$ in the final loss can constrain the degree of random sampling.
${{\mathbf{t}}}^{\text{sup}}={{\mathbf{t}}}+\frac{{{\mathbf{v}}}-{{\mathbf{t}}}}{\|{{\mathbf{v}}}-{{\mathbf{t}}}\|}\|\mathbf{r}\|.$
Appendix C Relational Graph Attention Network
In FRL (Sec. 3.2.1), for the text-frame graph with $n=1+S+M$ nodes, there are three types of edges (relations): text-text, text-frame, and frame-frame. So we have three adjacency matrices $\mathbf{A}^{r}\in\mathbb{R}^{n\times n}$ ($r\in\{tt,ff,tf\}$), and $\mathbf{A}_{ij}^{r}=1$ when node pair $(i,j)$ is linked by relation $r$. For example,
$\mathbf{A}_{ij}^{tt}=\left\{\begin{aligned} &1,&i,j\in[1,1+S],\\ &0,&\text{otherwise}.\end{aligned}\right$. (20)
Once the attention scores $\alpha_{ij}^{r,h}$ is calculated (Eq. 6), we can construct node attention matrix $\mathbf{E}^{r,h}\in\mathbb{R}^{n\times n}$ as
$\mathbf{E}_{ij}^{r,h}=\left\{\begin{aligned} &\alpha_{ij}^{r,h},&\mathbf{A}_{ij}^{r}=1,\\ &0,&\mathbf{A}_{ij}^{r}=0.\end{aligned}\right$. (21)
So the second term in Eq. 7 can be rewritten as
$\Big\|_{h=1}^{H}\sum_{r}\sum_{j\in\mathcal{N}_{i}^{r}}\alpha_{ij}^{r,h}\mathbf{W}^{r,h}\mathbf{h}_{j}$ (22)
$=$ $\Big\|_{h=1}^{H}\sum_{r}\sum_{j=1}^{n}\mathbf{E}_{ij}^{r,h}\mathbf{W}^{r,h}\mathbf{h}_{j},$
and it can be further written into matrix form as
$\Bigg[\Big\|_{h=1}^{H}\sum_{r}\mathbf{E}^{r,h}\mathbf{H}(\mathbf{W}^{r,h})^{\top}\Bigg]_{i},\\$ (23)
where $\mathbf{H}=[\mathbf{h}_{1}\|\mathbf{h}_{2}\|\cdots\|\mathbf{h}_{n}]^{\top}\in\mathbb{R}^{n\times d}$ is the node matrix of the intermedia layer. Before the first layer, $\mathbf{H}$ is the initial node matrix $\mathbf{X}$. And $\cdot]_{i}$” denotes the $i$-th row of the matrix. Finally, the output node representation (Eq. 7) can be rewritten into matrix form as
$\mathbf{H}^{\prime}=\sigma\Bigg(\mathbf{H}(\mathbf{W}_{\text{out}})^{\top}+\Big\|_{h=1}^{H}\sum_{r}\mathbf{E}^{r,h}\mathbf{H}(\mathbf{W}^{r,h})^{\top}\Bigg).$ (24)
Then $\mathbf{H}^{\prime}$ is input of the next layer.
In the final layer, the concatenation operation $\|_{h=1}^{H}$” is replaced by an average function across attention heads to obtain a $d$-dimentional output, and $\mathbf{W}_{\text{out}}\in\mathbb{R}^{d\times d}$.
Appendix D Training Energy-based Model
D.1 Derivations
Given a distribution $p(x)$, energy-based model (EBM) models it by Boltzmann distribution $p_{\theta}(x)=\frac{\exp(-E_{\theta}(x))}{Z_{\theta}}$. Its negative log-likelihood (NLL) loss is
$\mathcal{L}_{\theta}=-\mathbb{E}_{x\sim p(x)}[\log p_{\theta}(x)].$
To minimize $\mathcal{L}_{\theta}$, we need to calculate the gradient of $\log p_{\theta}(x)$ w.r.t. $\theta$:
$\nabla_{\theta}\log p_{\theta}(x)$
$=$ $\nabla_{\theta}\log\frac{\exp(-E_{\theta}(x))}{Z_{\theta}}$
$=$ $-\nabla_{\theta}E_{\theta}(x)-\nabla_{\theta}\log Z_{\theta}$
$=$ $-\nabla_{\theta}E_{\theta}(x)-\nabla_{\theta}\log\sum_{x}\exp(-E_{\theta}(x))$
$=$ $-\nabla_{\theta}E_{\theta}(x)-\frac{1}{\sum_{x^{\prime}}\exp(-E_{\theta}(x^{\prime}))}\sum_{x}\nabla_{\theta}\exp(-E_{\theta}(x))$
$=$ $-\nabla_{\theta}E_{\theta}(x)+\sum_{x}\frac{\exp(-E_{\theta}(x))}{\sum_{x^{\prime}}\exp(-E_{\theta}(x^{\prime}))}\nabla_{\theta}E_{\theta}(x)$
$=$ $-\nabla_{\theta}E_{\theta}(x)+\sum_{x}p_{\theta}(x)\nabla_{\theta}E_{\theta}(x)$
[…]
So we have
$\nabla_{\theta}\mathcal{L}_{\theta}$ $=-\nabla_{\theta}\mathbb{E}_{x\sim p(x)}[\log p_{\theta}(x)]$
[…]
Therefore, the EBM loss can be reformulated as
[…]
$\approx\frac{1}{B}\sum_{i=1}^{B}E_{\theta}(x_{i})-\frac{1}{B^{\prime}}\sum_{j=1}^{B^{\prime}}E_{\theta}(\tilde{x}_{j}),$
where $\{\tilde{x}_{j}\}_{j=1}^{B^{\prime}}$ is $B^{\prime}$ fake data sampled from $p_{\theta}(x)$. In practice, we set $B^{\prime}=B$ for convenience.
D.2 Sampling from EBM
Given a Boltzmann distribution $p_{\theta}(x)=\frac{\exp(-E_{\theta}(x))}{Z_{\theta}}$, the gradient of $\log p_{\theta}(x)$ w.r.t. $x$ is
$\nabla_{x}\log p_{\theta}(x)$ $=\nabla_{x}\log\frac{\exp(-E_{\theta}(x))}{Z_{\theta}}$
$=-\nabla_{x}E_{\theta}(x)-\nabla_{x}\log Z_{\theta}$
$=-\nabla_{x}E_{\theta}(x).$
So we can first initialize a sample using uniform distribution $x^{(0)}\sim\mathcal{U}$, and utilize multi-step stochastic gradient ascending to make the sample follows the distribution $p_{\theta}(x)$:
$x^{(k+1)}$ $=x^{(k)}+\eta\nabla_{x^{(k)}}\log p_{\theta}\left(x^{(k)}\right)+\epsilon^{(k)}$
$=x^{(k)}-\lambda\nabla_{x^{(k)}}E_{\theta}\left(x^{(k)}\right)+\epsilon^{(k)},$
where $\epsilon\sim\mathcal{N}\left(0,\sigma^{2}\right)$. After $K$ steps, $x^{(K)}$ is the output.
D.3 Replay Buffer
Furthermore, we use sample relay buffer [11]. The initialization of MCMC chain plays a crucial role in mixing time, but Langevin dynamics does not place restrictions on sample initialization given sufficient sampling steps. Thus, we use a sample replay buffer $\mathcal{R}$ in which we preserve previously generated samples and use either these samples or uniform noise for initialization.
D.4 Regularization for EBM Loss
[…]
$\mathcal{L}_{\text{ebm}}=$ $\frac{1}{B}\sum_{i=1}^{B}E_{\theta}(x_{i})-\frac{1}{B^{\prime}}\sum_{j=1}^{B^{\prime}}E_{\theta}(\tilde{x}_{j})$
$+c\Bigg[\frac{1}{B}\sum_{i=1}^{B}E_{\theta}(x_{i})^{2}+\frac{1}{B^{\prime}}\sum_{j=1}^{B^{\prime}}E_{\theta}(\tilde{x}_{j})^{2}\Bigg],$
where $c$ is a coefficient. We set $c=1.0$.
D.5 Algorithm for Sampling Text-Frame Pairs
To calculate the second term in energy loss (Eq. 15), a $K$-step MCMC sampling is conducted to sample $B^{\prime}$ text-frame pairs using energy function $E_{\theta}$. At the beginning of training, the replay buffers for text and frames are initialized as empty set:
$\mathcal{R}_{t}\leftarrow\varnothing,\ \ \ \mathcal{R}_{F}\leftarrow\varnothing,$
and the whole sampling algorithm is shown in Algorithm 1.
Algorithm 1 Sampling Algorithm of EAM.
Input: Energy function $E_{\theta}(\cdot,\cdot)$, replay buffers $\mathcal{R}_{t}$ and $\mathcal{R}_{F}$.
Parameter: MCMC step size $\eta$, MCMC noise variance $\sigma^{2}$, MCMC steps $K$, number of MCMC samples $B^{\prime}$.
Output: Sampled text-frame embedding pairs $\{\tilde{{{\mathbf{t}}}}_{j},\tilde{{\mathbf{F}}}_{j}\}_{i=1}^{B^{\prime}}$, and the updated replay buffers $\mathcal{R}_{t}$ and $\mathcal{R}_{F}$.
1:  for $j=1$ to $B^{\prime}$ do
2:   $\tilde{{{\mathbf{t}}}}_{j}^{(0)}\sim\mathcal{R}_{t}$ with $95\%$ probability and $\mathcal{U}$ otherwise;
3:   $\tilde{{\mathbf{F}}}_{j}^{(0)}\sim\mathcal{R}_{F}$ with $95\%$ probability and $\mathcal{U}$ otherwise;
4:   for $\text{sample step }k=1$ to $K$ do
5:    $\bm{\epsilon}_{t_{j}}^{(k)}\sim\mathcal{N}(0,\sigma^{2})$;
6:    $\tilde{{{\mathbf{t}}}}_{j}^{(k)}=\tilde{{{\mathbf{t}}}}_{j}^{(k-1)}-\eta\nabla_{\tilde{{{\mathbf{t}}}}_{j}^{(k-1)}}E_{\theta}(\tilde{{{\mathbf{t}}}}_{j}^{(k-1)},\tilde{{\mathbf{F}}}_{j}^{(k-1)})+\bm{\epsilon}_{t_{j}}^{(k)}$;
7:    $\bm{\epsilon}_{v_{j}}^{(k)}\sim\mathcal{N}(0,\sigma^{2})$;
8:    $\tilde{{\mathbf{F}}}_{j}^{(k)}=\tilde{{\mathbf{F}}}_{j}^{(k-1)}-\eta\nabla_{\tilde{{\mathbf{F}}}_{j}^{(k-1)}}E_{\theta}(\tilde{{{\mathbf{t}}}}_{j}^{(k-1)},\tilde{{\mathbf{F}}}_{j}^{(k-1)})+\bm{\epsilon}_{v_{j}}^{(k)}$;
9:   end for
10:   $\tilde{{{\mathbf{t}}}}_{j}\leftarrow\texttt{stop-gradient}\big[\tilde{{{\mathbf{t}}}}_{j}^{(K)}\big]$;
11:   $\mathcal{R}_{t}\leftarrow\mathcal{R}_{t}\cup\{\tilde{{{\mathbf{t}}}}_{j}\}$;
12:   $\tilde{{\mathbf{F}}}_{j}\leftarrow\texttt{stop-gradient}\big[\tilde{{\mathbf{F}}}_{j}^{(K)}\big]$;
13:   $\mathcal{R}_{F}\leftarrow\mathcal{R}_{F}\cup\{\tilde{{\mathbf{F}}}_{j}\}$;
14:  end for
15:  Sampled text-frame embedding pairs $\{\tilde{{{\mathbf{t}}}}_{j},\tilde{{\mathbf{F}}}_{j}\}_{i=1}^{B^{\prime}}$;
16:  Updated replay buffers $\mathcal{R}_{t}$ and $\mathcal{R}_{F}$.
Appendix E Experiment Settings
E.1 Datasets
We evaluate the proposed EagleNet using four widely used datasets.
(1) MSRVTT [52] includes 10,000 video clips, each of which has 20 captions. We employ the 1k-A split [31, 14] to use 9k videos for training and 1k videos for validation and test.
(2) DiDeMo [1] contains 10,642 Flickr video clips and 40,543 captions. We follow [28] to concatenate all captions of a video as a query.
(3) MSVD [4] is composed of 1,970 video clips with a length ranging from 1 to 62 seconds, each of which contains about 40 captions. We split the training, validation, and test sets with 1,200, 100, and 670 videos following [33].
(4) VATEX [45] consists of 34,991 video clips, each of which is related to multiple captions. We apply the same dataset split as [5] to select 25,911, 1,500, and 1,500 videos as training, validation, and test sets.
E.2 Evaluation Metrics
Following [33], we use the standard TVR metrics, including Recall (%) at rank 1/5/10 (R@1, R@5, R@10), Median Rank (MdR), Mean Rank (MnR), and Rsum (R@1++R@5++R@10). We also evaluate the average training time per epoch (sec./batch) and the total test time (sec.) on MSRVTT with ViT-B/16.
E.3 Selected baselines
We select three types of models as baselines for comparison. The first type is early works that are based on pre-extracted text and video features (not end-to-end models), including CE [31], MMT [14], HiT [29], and TeachText [9]. The second group is the end-to-end models that use pre-trained networks (e.g., ResNet [20] and CLIP) as encoders, including ClipBert [26], SupportSet [36], Fronzen [2], CLIP4Clip [33], TS2Net [32], CLIP-VIP [53], XPool [17], TMASS [43], RAP [3], TFVL [40], GLSCL [57], TempMe [41], TV-ProxyNet [51], and Video-ColBERT [38]. And our proposed EagleNet is an end-to-end model using pre-trained CLIP encoders as the backbone.
E.4 Implementation Details
The pre-trained CLIP model (ViT-B/32 and ViT-B/16) is used as image and text encoder [37]. We use CLIP4Clip [33] as backbone, and apply similar configurations, such that the hidden dimension $d=512$, weight decay $0.2$, dropout $0.3$, learning rate $10^{-7}$ for CLIP modules and $10^{-4}$ for non-CLIP modules; and we use a cosine scheduler for warmup, with the proportion of $0.1$. We uniformly sample $M=12$ frames for each video clip on all datasets ($64$ for DiDeMo). All frames are resized into $224\times 224$. We set the max length of text to $32$ for all datasets ($64$ for DiDeMo). For FRL, the number of text candidates $S=20$, the number of attention heads in RGAT $H=4$, and the number of RGAT layers $L=2$. For MCMC sampling in EAM, step number $K=20$, step size $\eta=1$, and noise variance $\sigma^{2}=0.005$. The sigmoid loss is initialized as $\tau_{p}=4.77$ and b=12.93b=-12.93 following [56]. For loss coefficient, we set $\lambda_{\text{sup}}=0.8$ following [43], and $\lambda_{\text{eam}}=1.0$ by default. We train all models on two A100 GPUs for $5$ epochs with a total batch size of $B=64$. In inference, we run on one A100 GPU with a batch size of $8$.
Appendix F Additional Analysis
Model Training Test
CLIP4Clip 2947.4 49.7
XPool 3062.7 51.0
TS2Net 4367.1 99.7
CLIP-VIP 2497.0 41.1
UCoFiA (t2v) 21021.4 2969.8
(v2t) 10000.3
ProST 3199.1 51.4
TMASS 3066.1 115.3
GLSCL 4078.2 148.5
TempMe 5271.7 18.2
TV-ProxyNet 2619.1 74.2
Video-ColBERT 3542.3 44.0
EagleNet (Ours) 5207.0 141.7
Table 6: Average training time per epoch and test time (sec.) consumption of different models (ViT-B/16) on MSRVTT.
F.1 Complexity Analysis
The main differences between our proposed EagleNet and other baselines lie in three components in the text-video similarity computation stage: (1) stochastic text embeddings sampling; (2) Fine-Grained Relationship Learning (FRL); (3) Energy-Aware Matching (EAM).
Stochastic Text Embeddings Sampling.
Following TMASS [43], we first calculate text-frame similarity and use a linear layer to determine the radius, so the complexity is given by $\mathcal{O}(B^{2}M^{2}d+B^{2}Md)=\mathcal{O}(B^{2}M^{2}d)$, where $B$ is batch size, $M$ is number of frames, and $d$ is the number of latent dimensions. Since we need to consider each text-video pair during the similarity computation, so the number of pairs is $B^{2}$. Then we use this radius to sample $S$ text embeddings, which adds a complexity of $\mathcal{O}(B^{2}Sd)$. The total complexity remains $\mathcal{O}(B^{2}M^{2}d)$ since $M^{2}>>S$.
Fine-Grained Relationship Learning.
This process primarily uses a $L$-layer $H$-head RGAT to process text-frames graph, which contain $n=1+S+M$ nodes. In each layer, RGAT first transforms the inputs, calculates the attention matrix, and obtain the output node features(Eqs. 5, 6 and 7), yielding a complexity of $\mathcal{O}(B^{2}RHnd^{2}+B^{2}RHnd+B^{2}RHn^{2}d)=\mathcal{O}(B^{2}RHnd^{2})$, where $R$ is the number of relationship types. The final text embedding (Eqs. 8, 9 and 10) is computed with a complexity of $\mathcal{O}(B^{2}Sd)$. So the complexity of each RGAT layer is $\mathcal{O}(B^{2}RHnd^{2})$, such that the final complexity is $\mathcal{O}(B^{2}LRHnd^{2})$.
Energy-Aware Matching.
In this part, the $K$-step MCMC sampling incurs the most time cost, which we focus on primarily. In each sample step, we use a 2-layer MLP as the energy function to calculate the scalar energy of a text-video pair and to calculate the gradient (Eqs. 14 and 16), so the total complexity is $\mathcal{O}(K{B^{\prime}}Md{d^{\prime}}+K{B^{\prime}}M{d^{\prime}})=\mathcal{O}(K{B^{\prime}}Md{d^{\prime}})$, where $B^{\prime}$} is the number of sampled text-video pairs (normally we set ${B^{\prime}}=B$), and $d^{\prime}$} is the number of latent dimensions of energy function. Since we consider the energy of each text-video pair, the number of samples is $B$. So the total complexity of EAM is $\mathcal{O}(KBMd{d^{\prime}})$.
In the training phase, the overall additional complexity compared to other baselines is the summation of all three components above, which is $\mathcal{O}(B^{2}LRHnd^{2})$. Since EAM is only conducted during training, the complexity for the test phase is the combination of the first and second components, which also results in $\mathcal{O}(B^{2}LRHnd^{2})$. Overall, the additional computational overhead of EagleNet primarily comes from FRL and EAM. This remains manageable in practice because $B$, $L$, $R$, $H$, and $n$ are typically small, and $d$ is a fixed constant. This indicates that our model incurs only modest additional runtime compared to baselines (Tab. 6).
F.2 Simple Alternative for Graph Encoder
As illustrated in Sec. 3.2.1, the constructed text-frame graph is a heterogeneous graph with multiple types of nodes and edges, we proposed to use RGAT (Eq. 7) for learning the edge attention scores. If we ignore the types of links and regard the text-frame graph as a fully connected graph, a simple GAT can be used as an alternative.
Specifically, we simply treat the text-frame graph as a fully-connected graph, i.e., $\mathbf{A}_{ij}\equiv 1$, so the edge weight is given by
$e_{ij}^{h}$ $=\psi\bigg(\left[\mathbf{W}^{h}\mathbf{h}_{i}\|\mathbf{W}^{h}\mathbf{h}_{j}\right]\bigg)\in\mathbb{R},$ (25)
then the attention score is calculated as
$\alpha_{ij}^{h}$ $=\frac{\exp\left(\operatorname{LeakyReLU}\left(e_{ij}^{h}\right)\right)}{\sum_{k=1}^{n}\exp\left(\operatorname{LeakyReLU}\left(e_{ik}^{h}\right)\right)},$ (26)
where $n=1+S+M$ is the number of nodes, and we can construct the node attention matrix $\mathbf{E}^{h}=\{\alpha_{ij}^{h}\}_{ij}\in\mathbb{R}^{n\times n}$.
The output node matrix of each layer is
$\mathbf{H}^{\prime}=\sigma\left(\Big\|_{h=1}^{H}\mathbf{E}^{h}\mathbf{H}(\mathbf{W}^{h})^{\top}\right).$ (27)
In the final layer, the concatenation operation $\|_{h=1}^{H}$” is replaced by an average function following Appendix C.
Similarly, we select the edge weights from the final layer and average them across attention heads to get the final text-frame weights, i.e.,
$e_{ij}=\frac{1}{H}\sum_{h=1}^{H}e_{ij}^{h},\ \ \begin{aligned} &i\in[1,1+S],\\ &j\in[1+S+1,1+S+M],\end{aligned}$ (28)
Graph MSRVTT DiDeMo
Encoder R@1↑ R@5↑ R@10↑ R@1↑ R@5↑ R@10↑
w/o EAM
GAT 49.4 76.8 85.9 47.5 76.6 85.9
RGAT 49.9 77.3 84.3 49.1 77.5 85.0
w/ EAM
GAT 49.9 76.6 85.2 48.9 78.4 86.6
RGAT 51.0 76.2 85.6 51.5 78.7 86.8
Table 7: Text-to-video on different graph encoders.
From Tab. 7, Overall, RGAT performs better than GAT on both MSRVTT and DiDeMo with or without EAM. This illustrates that RGAT is better at learning multi-relationships than GAT. Moreover, the performance will improve when employing EAM, further demonstrating the effectiveness of our proposed EAM in enhancing FRL and the final performance. We select RGAT as the graph encoder in the final implementation.
F.3 Ablation Study for Coefficient of Loss Function
For the loss coefficient, the setting of $\lambda_{sup}=0.8$ follows the practice in TMASS [43], and the coefficient $\lambda_{eam}$ was set to 1 by default to maintain a stable and consistent gradient contribution from EAM. We further conduct a sensitivity analysis on the loss coefficients. Since $\lambda_{sup}=0.8$ has been validated as an effective choice for balancing supervision signals for balancing supervision signals [43], we focus our analysis on $\lambda_{eam}$.
From Tab. 8, the performance remains stable across various $\lambda_{eam}$, and the best overall performance is achieved when $\lambda_{eam}=1.0$. This indicates that the proposed model is not sensitive to the choice of $\lambda_{eam}$, which supports our default setting for simplicity and robustness.
$\lambda_{eam}$ MSRVTT DiDeMo
R@1↑ R@5↑ R@10↑ R@1↑ R@5↑ R@10↑
0.1 50.7 76.6 85.7 50.9 78.9 86.6
0.5 50.1 76.5 86.2 51.2 78.6 86.6
1.0 51.0 76.2 85.6 51.5 78.7 86.8
Table 8: Text-to-video on different $\lambda_{eam}$.
F.4 Ablation Study for Frame-to-Frame Edges
To verify the effectiveness of modeling internal frame-to-frame relationships within a video, we evaluate a variant of EagleNet that removes only the frame-to-frame edges (f2f) in FRL.
As shown in Tab. 9, comparing the “w/o f2f” variant with our full EagleNet reveals a consistent performance improvement across both datasets (e.g., +0.4% R@1 on MSRVTT and +1.6% R@1 on DiDeMo). This gain directly validates the importance of modeling inter-frame dependencies, as it enables the RGAT to capture temporal continuity and the logical progression of video content, rather than treating frames as independent, isolated entities.
MSRVTT DiDeMo
R@1↑ R@5↑ R@10↑ R@1↑ R@5↑ R@10↑
w/o FRL 48.6 75.3 83.7 43.8 75.6 83.8
w/o f2f 50.6 77.5 84.6 49.9 78.7 86.5
EagleNet 51.0 76.2 85.6 51.5 78.7 86.8
Table 9: Text-to-video on frame-to-frame edges (f2f).
F.5 Visualization
F.5.1 Different Models
In Fig. 4 we visualize some text-to-video retrieval results by our proposed EagleNet, TMASS [43], and TV-ProxyNet [51].
TMASS and TV-ProxyNet do not exploit the internal frame-to-frame relationships and the video context, so they might retrieve incorrect videos with similar semantics but different contexts. For example, in Fig. 4 (A), TV-ProxyNet retrieves the video where there is a man and a car, ignoring that the man is driving the car but not giving a review. Similarly, in Fig. 4 (B), TMASS retrieves the video with Pokémon, but they are not dancing.
Furthermore, as previously mentioned, the enriched text embeddings generated by TMASS can negatively impact retrieval outcomes due to the inherent randomness from stochastic sampling. However, EagleNet mitigates this individual bias. As illustrated in  Fig. 4, EagleNet successfully retrieves the correct video clip, while TMASS retrieves an incorrect clip with similar semantics. For instance, given the query text “sports vine clips of football,” EagleNet retrieves the appropriate video, whereas TMASS provides a video showcasing various sports, including baseball and American football. This highlights the accuracy of our generated text embeddings, whereas TMASS introduces some undesirable information into the embeddings.
F.5.2 Different Components
In Fig. 5 we visualize results by different components, baseline, without FRL, without EAM, and our proposed EagleNet. In the absence of one component, the model might lose some key information, leading to a similar but incorrect video.
For example, the enriched text cannot capture frame contextual information when not without FRL (‘Base’ and ‘Ours w/o FRL’). In Fig. 5 (B), the “w/o FRL ” variant only captures the elements “speaks” and “children in a classroom” in the returned video, while ignoring the mismatch between the “woman” in the video and the “man” mentioned in the query. Similarly, in Fig. 5 (C), both “Base” and “w/o FRL” retrieve videos as long as one or two of the mentioned characters appear, without considering whether all the characters mentioned in the query are actually present throughout the video.
Figure 4: Visualization of text-to-video retrieval results by our proposed EagleNet, TMASS [43], and TV-ProxyNet [51]. Green denotes the correct retrieval result, and red is the wrong result.
Figure 5: Visualization of text-to-video retrieval results by different components of EagleNet. Green denotes the correct retrieval result, and red is the wrong result.